OpenStudy (anonymous):

A manufacturer of toys sells a certain no. of toys to a wholesale dealer at a profit of 25%. The wholesale dealer sells these toys to a retail merchant at a profit of 20%. If the retailer makes a profit of 50/3% by selling these toys for Rs 5250, find the manufacturer's cost of toys.

OpenStudy (anonymous):

can u solve it?

OpenStudy (anonymous):

Let the cost of production be x: Now total profit by the manufacturer, dealer and retailer: 25% of 20% of 50/3% of x = 5250 \[\large \frac{25}{100} \times \frac{20}{100} \times \frac{50}{3 \times 100} \times x = 5250\] Can you solve it now?

OpenStudy (anonymous):

50400

OpenStudy (anonymous):

Let me check..

OpenStudy (anonymous):

oh no its wrong. wait

OpenStudy (anonymous):

Check it again..

OpenStudy (anonymous):

it would be 630000

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Yes Right..

OpenStudy (anonymous):

thanx i took 1/4 instead of taking 1/5..